# Local-Cloud Inference Offloading for LLMs in Multi-Modal, Multi-Task, Multi-Dialogue Settings

### v1

@@ III-B Local-Cloud LLM Inference
where $R_{t}$, $P_{t}$, and $\mathcal{M}_{t}\subseteq\mathcal{M}$ represent the LLM response, text prompt, and the set of selected modalities uploaded to the cloud, respectively, at dialogue $t$. $D^{(m)}$ denotes the data source of modality $m\in\mathcal{M}$, while $\text{LLM}_{\text{Local}}$ and $\text{LLM}_{\text{Cloud}}$ indicate the local LLM and the cloud LLM, respectively. Thanks to the transformer architecture of LLM, it is capable of maintaining a contextual memory over extended dialogues, thereby allowing the LLM to generate responses for the current prompt based on the accumulated context of all previous prompts and responses. Hence, Eq. (1) can be rewritten to reflect the cumulative nature of the prompts. For example, in the case of the cloud LLM with multi-modal inputs,
$R_{t}=\text{LLM}_{\text{Cloud}}\left(P_{t},\bigcup_{i=0}^{t}\bigcup_{m\in\mathcal{M}_{i}}\{D^{(m)}\},\bigcup_{i=0}^{t-1}\{P_{i},R_{i}\}\right),$ (2)
@@
where $t=1,2,\dots,T$. In Eq. (2), in addition to the inputs in Eq. (1), we consider the previously selected modalities as well as prior prompts and responses from the series of earlier dialogues $i=0,1,\dots,t-1$.
@@
Objective of LCIO. The primary objective in deploying the LCIO system is to strategically decide (i) whether to use the local LLM or cloud LLM and (ii) which multi-modal data sources to utilize, for each task or dialogue. This decision-making process aims to enhance response quality while efficiently managing latency and costs throughout extended interactions. The goal is to ensure a seamless user experience in complex multi-dialogue environments, particularly under the constraints of available resources (e.g., latency and costs).
@@
Challenges in LCIO. Fig. 2 summarizes the key challenges of our problem. Selecting the appropriate LLM and modalities for different tasks remains a complex decision, particularly in each dialogue where the choice between utilizing cloud LLM or local LLM is influenced by the distinct difficulties of prompts and the availability of multi-modal data sources. More importantly, while humans may be capable of making short-term decisions for current tasks without history (for example, intuitively deciding to use the local LLM for simple message editing), the challenge intensifies when considering the entire spectrum of interactions over prolonged multi-dialogues. Here, we need to consider response quality, latency, and costs cumulatively throughout the multi-dialogue process. This includes decisions on offloading inference to the cloud, where users need to not only determine which and how many multi-modal data sources to transmit but also balance the quality of the response with other metrics, such as latency and cost. Furthermore, early uploading of modality data to the cloud LLM can facilitate a more quick and nuanced understanding of the environment, which subtly enhances the management of long-term, multi-dialogue interactions.
III-C RCRL for LLM and Modality Selection
We propose a resource-constrained RL (RCRL) approach to make informed decisions regarding the selection between local and cloud LLM services, as well as the determination of which data modalities to be uploaded, as shown in Fig. 3 and Alg. 1. In the following, we describe the details of our RL components.
Environment: The environment is defined by dialogues between the user and the LLM services.
@@ III-C RCRL for LLM and Modality Selection
State: The state captures both the history of previous LLM selections and the history of uploaded modalities, along with the current task category $n$. Formally, each state in the state space $\mathcal{S}$ can be visualized as a repeated sequence of the tuple $(\text{LLM},\mathcal{M},n)$, spanning $\tau$ time steps. Thus, we can represent the state space $\mathcal{S}$ as follows:
$\mathcal{S}=(\text{LLM},\mathcal{M},n)^{\tau},$ (3)
@@
where LLM indicates the choice between local LLM and cloud LLM. The superscript $\tau$ indicates that there are $\tau$ state components over $\tau$ consecutive dialogues, showing the history of model and modality selections.
@@
Action: The action space consists of the choice between utilizing local or cloud LLM services, as well as the selection of one or multiple data modalities for processing. Therefore, the action space $\mathcal{A}$ can be described as follows:
$\mathcal{A}=(\text{LLM},\mathcal{M}).$ (4)
Figure 3: Illustration of RCRL within the LCIO System. The reward computation step is only necessary during the training process. Once the RL policy is deployed, it can determine the appropriate action solely based on the state without processing any external information from the environment (e.g., prompts or responses from the LLM), enhancing the real-time responsibility of the RL action delivery.
@@
$r_{t}=\alpha S_{R_{t}}^{\prime}+\beta_{\Lambda}\sum_{m\in\mathcal{M}_{t}}\Lambda\left(P_{t},D^{(m)}\right)-\beta_{\psi}\psi_{a_{t}}-\beta_{\phi}\phi_{a_{t}},$ (5)
@@
  • Estimated Response Score $S_{R_{t}}^{\prime}$ assesses how effectively the LLM response $R_{t}$ (from either local or cloud) meets the requirements of the text prompt. The response score is used solely for RL training purposes and is not required during the inference phase. A more detailed description of this metric is provided in Sec. III-D.
@@
  • Association $\Lambda\left(P_{t},D^{(m)}\right)$ quantifies the relevance of the multi-modal data $D^{(m)}$ to the prompt $P_{t}$. We aggregate the association scores for the set of uploaded modalities $\mathcal{M}_{t}$, which depend on the current action $a_{t}$. Further details are provided in Sec. III-E.
@@
  • Latency $\psi_{a_{t}}$ represents the latency of the LLM inference under different actions $a_{t}$. This includes the computation time of the local LLM or the interaction time with the cloud LLM for uploading the selected data modalities. Refer to Sec. III-F for the formal definition.
@@
  • Usage Cost $\phi_{a_{t}}$ reflects the cost of LLM under different actions $a_{t}$, including the energy consumption of the local LLM or the service fee of the cloud LLM for various uploaded modalities, as described in Sec. III-G.
@@
The estimated response score $S_{R_{t}}^{\prime}$ and the association $\Lambda\left(P,D^{(m)}\right)$ are considered positive indicators (higher is better, $\uparrow$). In contrast, latency $\psi_{a_{t}}$ and usage cost $\phi_{a_{t}}$ are treated as negative indicators (lower is better, $\downarrow$). Variables $\alpha$, $\beta_{\Lambda}$, $\beta_{\psi}$, and $\beta_{\phi}$ are weighting factors that adjust the influence of each component on the total reward, allowing the system to prioritize different aspects according to specific operational goals and constraints.
@@
Markov Decision Process (MDP) and Optimization: Based on our state and action definitions, we can model this problem as an MDP, where the learning objective is to maximize the cumulative reward, similar to other RL works. In each dialogue $t$, RL algorithms observe the state $s_{t}$, select an action $a_{t}$, obtain a transition probability $P(s_{t+1}|s_{t},a_{t})$, receive a reward $r_{t}$, and transition to the next state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. Given the current state $s\in\mathcal{S}$, the action $a\in\mathcal{A}$ is selected according to a specific policy $\pi$ as $a\sim\pi(\cdot|s)$, where $\pi(a|s)$ represents the probability of selecting the action $a$ in state $s$. The value function $V^{\pi}(s)$ and the action-value function $Q^{\pi}(s,a)$, respectively, predict the expected discounted rewards of states and actions, following the policy $\pi$:
$V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\bigm{|}s_{0}=s\right],$ (6)
$Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\right].$ (7)
@@
The policy is parameterized by $\theta$ and the optimization of policy $\pi_{\theta}$ aims to maximize the expected discounted rewards:
$\max_{\theta}J(\theta)\triangleq\mathbb{E}_{s,a\sim\pi_{\theta}}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right].$ (8)
@@
In deep RL optimization, various strategies can be employed to adjust the parameters $\theta$ of the policy $\pi_{\theta}(a|s)$ with the goal of maximizing the performance metric $J(\theta)$. To achieve this, a loss function $f(\theta)$ is defined such that minimizing $f(\theta)$ leads to maximizing $J(\theta)$. Taking Advantage Actor Critic (A2C) [35] as an example, the loss function $f_{\text{A2C}}(\theta)$ includes both the policy loss and the value loss, along with an entropy term to encourage exploration:
$f_{\text{A2C}}(\theta)$ $=\mathbb{E}_{s,a\sim\pi_{\theta}}[-\log\pi_{\theta}(a|s)A(s,a)$ (9)
$+\frac{1}{2}(V^{\pi_{\theta}}(s)-r)^{2}-\zeta H(\pi_{\theta}(\cdot|s))],$
@@
where $A(s,a)=r+\gamma V^{\pi_{\theta}}(s^{\prime})-V^{\pi_{\theta}}(s)$ is an estimate of the advantage function, providing a measure of the relative value of taking action $a$ in state $s$. The coefficient $\zeta$ controls the weight of the entropy bonus $H(\pi_{\theta}(\cdot|s))$. In Sec. IV-B, we also consider Proximal Policy Optimization (PPO) [36] and Deep Q Network (DQN) [37] versions of Eq. (9).
Algorithm 1 Proposed LCIO System
@@
2:    Sample a mini-batch of transitions $\{(s,a,s^{\prime})\}$ and corresponding metrics $\{(\Lambda,S_{R},\psi_{a},\phi_{a})\}$ from $\mathcal{D}$. $\triangleright$ Alg. 2
@@
4:    Estimate uncertain response scores $S_{R}^{\prime}$ for pending state-action pairs $(s^{\prime},a^{\prime})$ via nearest neighbor. $\triangleright$ Eq. (12)
@@
Resource-Constrained RL (RCRL): A key challenge in non-resource-constrained RL is the unrestricted maximization of rewards, leading to excessive resource use. Even when penalties like latency and usage costs are included in the reward function, they may not enforce strict resource limits, risking over-utilization. To ensure adherence to resource constraints, we incorporate RCRL to enhance decision-making capabilities in environments with limited resources. RCRL involves not only maximizing cumulative rewards, but also satisfying certain constraints on resource consumption such as latency or usage cost. Formally, we embed resource constraints into MDP by introducing a resource consumption function $C(a)$, which quantifies the resources consumed when action $a$ is taken in state $s$. The objective of RCRL is to minimize the loss function $f(\theta)$ (e.g., Eq. (9) for the A2C loss) while ensuring compliance with resource constraints, by solving
$\min_{\theta}f(\theta)\quad\mathrm{s.t.}\quad\sum_{i=t}^{t+\tau}C_{j}(a_{i})\leq\xi_{j},\;\text{for all}\;j\in\mathcal{J},$ (10)
@@
where $C_{j}(a)$ denotes the consumption of the $j$-th constrained resource with $\xi_{j}$ as the corresponding threshold, and $\mathcal{J}$ is the set that contains all constraints. $\tau$ is a time span, defined in Eq. (3). We transform resource constraints into a policy optimization problem using a penalty term to enforce constraints. Specifically, we define $g_{j}(\theta)\triangleq\sum_{i=t}^{t+\tau}C_{j}(s_{i},a_{i})-\xi_{j}$ to penalize constraint violations and aim to minimize the following overall loss $F(\theta)$:
$F(\theta)=f(\theta)+\lambda\sum_{j}\max\left\{g_{j}(\theta),0\right\},$ (11)
@@
where $\lambda$ denotes the penalty coefficient associated with the constraints. In our experiments, we consider two constraint penalty terms $g_{1}(\theta)$ and $g_{2}(\theta)$ on latency and usage cost, respectively.
III-D Uncertain Response Score Estimation
@@
Due to the enormous number of parameters in LLM, most LLM-related training strategies, such as RLHF and similar approaches, adopt offline training to mitigate substantial inference costs [38, 39]. Similarly, the RCRL framework in our LCIO system also requires learning and exploration of LLM performance in multi-modal, multi-task, and multi-dialogue settings, making a dedicated dataset essential for training RCRL. Given the nature of multi-task settings, particularly in generative tasks, there are often no explicit evaluation metrics, such as accuracy, task success rate, or error rate. This creates a significant challenge in assessing the quality of LLM responses. Moreover, human annotations are not only costly but also inconsistent, individuals evaluating the same LLM response may provide differing assessments, and these variations become even more pronounced across different evaluators, further complicating the evaluation of response quality. In addition, the exploratory nature of RL introduces scenarios, where specific state-action pairs may not exist in the dataset, posing further challenges for quality assessment.
@@
To address these issues, we propose using LLM-as-judge to evaluate response quality. Following this, we employ a nearest neighbor strategy to estimate uncertain response scores, providing a method for handling the inherent uncertainties in outputs of local/cloud LLMs and the judge LLM.

### v3

@@ 3.2. Local-Cloud LLM Inference
where $R_{t}$ represents the response to the particular combination of LLM and modalities provided for dialogue $t$. $D^{(m)}$ denotes the data source of modality $m\in\mathcal{M}_{t}$, while $\text{LLM}_{\text{Local}}(\cdot)$ and $\text{LLM}_{\text{Cloud}}(\cdot)$ are representations of the local LLM and the cloud LLM in functional form. The transformer architecture of LLM is known for its capability of maintaining a contextual memory over extended dialogues, thereby allowing the LLM to generate responses for the current prompt based on the accumulated context of all previous prompts and responses, though the quality of the responses will differ between local and cloud LLMs. Hence, Eq. (1) can be rewritten to reflect the cumulative nature of the prompts. For example, in the case of the cloud LLM with multi-modal inputs,
(2) $\small R_{t}=\text{LLM}_{\text{Cloud}}\left(P_{t},\;\;\bigcup_{i=0}^{t}\bigcup_{m\in\mathcal{M}_{i}}\{D^{(m)}\},\;\;\bigcup_{i=0}^{t-1}\{P_{i},R_{i}\}\right),$
@@
where $t=1,2,\dots,T$. This further shows the dependency on the previously selected modalities as well as prior prompts and responses from the series of earlier dialogues $i=0,1,\dots,t-1$.
@@
TMO Objective and Rationale. The primary objective of TMO is to strategically decide (i) whether to use the local LLM or cloud LLM and (ii) which multi-modal data sources to utilize, for each task and dialogue. This decision-making process aims to enhance response quality while efficiently managing latency and costs throughout extended interactions. The goal is to ensure high response quality in complex multi-dialogue environments, particularly under the system constraints (e.g., latency and costs).
@@
The complexity of LLM and modality selection can be illustrated by the human decision-making process in such a setting. For example, iPhone users encounter choices among multiple LLMs (e.g., local Siri and the ChatGPT application), while also navigating various multi-modal data sources on mobile devices (e.g., front camera, rear camera, and screenshots). Although users can make short-term decisions for current tasks (for example, intuitively deciding to use the local LLM for simple message editing), the challenge intensifies when considering the entire spectrum of interactions over prolonged multi-dialogues and evolving contexts. Individuals must determine which of many multi-modal data sources to transmit and balance the quality of the response with other metrics, such as latency and various costs. Furthermore, early uploading of modality data to the cloud LLM can facilitate quicker and more nuanced understanding of the environment, enhancing long-term multi-dialogue interaction management. Optimal decision-making in these multi-modal, multi-task, and multi-dialogue scenarios is thus difficult by human inspection.
3.3. LLM and Modality Selection
We propose a resource-constrained RL (RCRL) approach to orchestrate the TMO system. Its key components are shown in Fig. 2, and Alg. 1 summarizes the process.
3.3.1. RL Formulation
We first formulate different components of the RL policy as follows.
@@ 3.3.1. RL Formulation
State: The state captures both the history of previous LLM selections and the history of uploaded modalities, along with the current inference task category $n\in\{1,...,N\}$. Formally, each state in the state space $\mathcal{S}$ can be visualized as a repeated sequence of the tuple $(\ell,\mathcal{M}^{\ell},n)$. Thus, we can represent the state space $\mathcal{S}$ over $\tau$ consecutive dialogues as follows:
(3) $\mathcal{S}=(\ell,\mathcal{M}^{\ell},n)^{\tau},$
@@
where $\ell\in\{\text{LLM}_{\text{Local}},\text{LLM}_{\text{Cloud}}\}$ indicates the choice between local LLM and cloud LLM. $\mathcal{M}^{\ell}=\mathcal{M}$ if $\ell=\text{LLM}_{\text{Cloud}}$ and $\mathcal{M}^{\ell}=\emptyset$ otherwise. (We will often refer to $\mathcal{M}^{\ell}$ as simply $\mathcal{M}$ when the context is clear.) The superscript $\tau$ indicates that there are $\tau$ state components over $\tau$ consecutive dialogues, showing the history of model and modality selections. The state $s_{t}\in\mathcal{S}$ for dialogue $t$ is then represented by the tuple $s_{t}=(\ell,\mathcal{M},n)_{t}$.
@@
Action: The action taken for the next dialogue consists of the choice between utilizing local or cloud LLM services, as well as the selection of one or multiple data modalities for inference processing. Therefore, the action space $\mathcal{A}$ can be described as follows:
(4) $\mathcal{A}=(\ell,\mathcal{M}^{\ell}),$
with the action $a_{t}=(\ell,\mathcal{M}^{\ell})_{t}$ for dialogue $t$ consisting of a choice of LLM and modalities to upload.
Figure 2. Illustration of RCRL within the TMO System.
@@
(5) $\small r_{t}=\alpha S_{R_{t}}+\beta_{\Lambda}\sum_{m\in\mathcal{M}_{t}}\Lambda\left(P_{t},D^{(m)}\right)-\beta_{\psi}\psi_{a_{t}}-\beta_{\phi}\phi_{a_{t}},$
@@
  • The response score $S_{R_{t}}$ assesses how effectively the LLM response $R_{t}$ (from either local or cloud) meets the requirements of the text prompt. The response score is used solely for RL training purposes and is not required during the inference phase. A more detailed description of this metric and how we estimate it is provided in Sec. 3.4.
@@
  • The association score $\Lambda(P_{t},D^{(m)})$ quantifies the relevance of multi-modal data $D^{(m)}$ to prompt $P_{t}$. We aggregate the association scores for the uploaded modality set $\mathcal{M}_{t}$, which depend on the current action $a_{t}$. Further details are provided in Sec. 3.5.
@@
  • The latency $\psi_{a_{t}}$ represents the latency of the LLM inference under different actions $a_{t}$. This includes the computation time of the local LLM or the interaction time with the cloud LLM for uploading the selected data modalities. Refer to Sec. 3.6 for the formal definition.
@@
  • The usage cost $\phi_{a_{t}}$ reflects the cost incurred by the LLM under action $a_{t}$, including the energy consumption of the local LLM or the service fee of the cloud LLM for various uploaded modalities, as described in Sec. 3.7.
@@
The response score $S_{R_{t}}$ and the association $\Lambda(P,D^{(m)})$ are thus positive indicators (higher is better, $\uparrow$). In contrast, latency $\psi_{a_{t}}$ and usage cost $\phi_{a_{t}}$ are negative indicators (lower is better, $\downarrow$). Variables $\alpha$, $\beta_{\Lambda}$, $\beta_{\psi}$, and $\beta_{\phi}$ are weighting factors that adjust the influence of each component on the total reward.
3.3.2. MDP and Optimization
@@
Based on our state and action definitions, we model this problem as a markov decision process (MDP), where the learning objective is to maximize the cumulative reward. In each dialogue $t$, RL algorithms observe the state $s_{t}$, select an action $a_{t}$, obtain a transition probability $P(s_{t+1}|s_{t},a_{t})$, receive a reward $r_{t}$, and transition to the next state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. Given the current state $s_{t}$, the action $a_{t}$ is selected according to a specific policy $\pi$ as $a\sim\pi(\cdot|s)$, where $\pi(a|s)$ represents the probability of selecting action $a$ in state $s$. The value function $V^{\pi}(s)$ predicts the expected discounted rewards of states following the policy $\pi$:
(6) $\small V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\bigm{|}s_{0}=s\right]$
@@
with discount $\gamma$. The optimization of policy $\pi_{\theta}$ with respect to its parameterization $\theta$ aims to maximize the expected discounted rewards in LLM inference offloading and modality selection:
(7) $\small\max_{\theta}J(\theta)\triangleq\mathbb{E}_{s,a\sim\pi_{\theta}}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right].$
@@
In deep RL optimization, various strategies can be employed to adjust the parameters $\theta$ of the policy $\pi_{\theta}(a|s)$ with the goal of maximizing $J(\theta)$. To achieve this, a loss function $f(\theta)$ is defined such that minimizing $f(\theta)$ leads to maximizing $J(\theta)$. Taking Advantage Actor Critic (A2C) (Mnih et al., 2016) as an example, the loss function $f_{\text{A2C}}(\theta)$ includes both the policy loss and the value loss, along with an entropy term to encourage exploration:
$f_{\text{A2C}}(\theta)=\mathbb{E}_{s,a\sim\pi_{\theta}}[-\log\pi_{\theta}(a|s)A(s,a)+\frac{1}{2}(V^{\pi_{\theta}}(s)-r)^{2}-\zeta H(\pi_{\theta}(\cdot|s))],$
@@
where $A(s,a)=r+\gamma V^{\pi_{\theta}}(s^{\prime})-V^{\pi_{\theta}}(s)$ for $s^{\prime}\neq s$ is an estimate of the advantage function, providing a measure of the relative value of taking action $a$ in state $s$. The coefficient $\zeta$ controls the weight of the entropy bonus $H(\pi_{\theta}(\cdot|s))$. In Sec. 4.2, we also consider Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Deep Q Network (DQN) (Mnih et al., 2015) versions of Eq. (8).
Algorithm 1 Local-Cloud Inference Offloading in TMO
@@
2:  Sample a mini-batch of transitions $\{(s,a,s^{\prime})\}$ and corresponding metrics $\{(\Lambda,S_{R},\psi_{a},\phi_{a})\}$ from $\mathcal{D}$.
@@
4:  Estimate uncertain response scores $S_{R}^{\prime}$ for pending state-action pairs $(s^{\prime},a^{\prime})$ via nearest neighbors. $\triangleright$ Eq. (11)
@@ 3.3.3. Resource-Constrained RL
We consider RCRL to enhance decision-making capabilities for TMO in environments with limited resources. RCRL involves not only maximizing cumulative rewards, but also satisfying certain constraints on resource consumption such as latency or usage cost incurred from processing different modalities. Formally, we consider a resource consumption function $C_{j}(a)$, which quantifies the amount of resource $j\in\mathcal{J}$ consumed when action $a$ is taken. RCRL then aims to minimize the loss function $f(\theta)$ (e.g., Eq. (8) for the A2C loss) while ensuring compliance with resource constraints, by solving
(9) $\min_{\theta}f(\theta)\quad\mathrm{s.t.}\quad\sum_{i=t}^{t+\tau}C_{j}(a_{i})\leq\xi_{j},\;\text{for all}\;j\in\mathcal{J},$
@@
where $C_{j}(a)$ denotes the consumption of the $j$-th constrained resource with $\xi_{j}$ as the corresponding budget over the horizon, and $\mathcal{J}$ is the set of all constraints. $\tau$ is the planning horizon length, defined in Eq. (3). We transform resource constraints into a policy optimization problem using a penalty term to enforce constraints. Specifically, we define $g_{j}(\theta)\triangleq\sum_{i=t}^{t+\tau}C_{j}(a_{i})-\xi_{j}$ to penalize constraint violations and aim to minimize the overall loss $F(\theta)$:
(10) $F(\theta)=f(\theta)+\lambda\sum_{j\in\mathcal{J}}\max\left\{g_{j}(\theta),0\right\},$
@@
where $\lambda\geq 0$ denotes the penalty associated with the constraints. In our experiments, we will consider two constraint penalty terms $g_{1}(\theta)$ and $g_{2}(\theta)$ on latency and usage cost, respectively.
3.4. Response Score Uncertainty Estimation
@@
Due to the vast number of parameters in LLMs, offline training approaches have become essential to mitigate substantial inference costs and enable future reusability (Ouyang et al., 2022). TMO’s RCRL procedure requires a specialized dataset for learning and exploration of LLM performance in multi-modal, multi-task, and multi-dialogue settings. Moreover, given the nature of multi-task settings, particularly in generative tasks, there often exist no explicit evaluation metrics, such as accuracy, task success rate, or error rate. Furthermore, the exploratory nature of RL, which frequently encounters new state-action pairs, presents significant evaluation challenges.
@@
To address these issues, we propose leveraging the notion of LLM-as-Judge to evaluate response quality. Following this approach, we employ a nearest neighbor strategy to estimate uncertain response scores, providing a method for handling the inherent uncertainties in four LLMs: the outputs of the local LLM, the cloud LLM, a reference LLM, and a judge LLM.
3.4.1. Response Quality Evaluation